# MMLongBench: Benchmarking Long-Context Vision-Language Models Effectively and Thoroughly

### v2

@@ 1 Introduction
To support such capabilities, researchers have proposed various techniques to extend the context windows of LVLMs, as seen in models such as LongVILA [9] and GPT-4o [10]. However, the development of effective evaluation benchmarks is lagging behind. It remains unclear how well current LCVLMs perform in long-context settings, what types of tasks they struggle with, and how robust they are to input length variation. Here, we take a closer look and find that existing benchmarks suffer from the following shortcomings and provide a summary of key feature comparisons in Table 1:
@@
Table 1: Comparison of benchmarks for LCVLMs: MM-NIAH [18], Visual Haystack [16], MMNeedle [11], MMLongBench-Doc (MMLB-Doc) [5], M-Longdoc [21], LongDocURL [17], and our MMLongBench. “Summ” and “DocVQA” refer to summarization and long-document visual question answering (VQA). “Mixed” indicates both natural and synthetic images. “$L$” denotes the number of input tokens, and “$L$ Control” means counting both text and image tokens together.
@@ 2 Related Work
In contrast, the evaluation of LVLMs’ long-context capability remains limited. Existing benchmarks only involve either NIAH [18, 11, 16, 65] or long-document VQA [5, 17], lacking comprehensive coverage across diverse vision-language applications. As a result, frontier LCVLMs [10, 33] only report long-context performance on other modalities, such as video [9, 66, 67] or audio [33, 10], and neglect the prevalent use cases of long-context vision-language inputs. While recent MileBench [68] claims to be a comprehensive long-context benchmark with various text-image tasks, our closer inspection reveals that it actually contains a lot of short-context tasks, and the average length is only about 9K tokens. Datasets like DocVQA [69], WebQA [70], and OCRVQA [71] contain only one image per sample and minimal context, making MileBench unqualified as a true long-context benchmark. In this work, we introduce the first comprehensive benchmark that evaluates a wide range of vision-language downstream applications across five standardized input lengths.
@@ 3 Our Benchmark — MMLongBench
In our work, we seek to address the limitations of current benchmarks by meeting the following criteria: i) broad coverage of both diverse vision-language downstream tasks and different image types, ii) a unified token-count method across different modalities and datasets, and iii) multiple standardized context lengths for each example, ranging from 8K to 128K tokens. In this section, we describe the task categories and datasets included in MMLongBench, highlighting how they address the limitations of existing evaluation benchmarks. An overview of MMLongBench is provided in Table 2, and several concrete examples are given in Appendix E.
@@
MM-NIAH-Ret SubEM/Acc Mixed 1,200 Retrieve text/image needles in web pages
MM-NIAH-Count Acc Mixed 1,178 Count text/image needles in web pages
MM-NIAH-Reason SubEM/Acc Mixed 1,158 Reason about text/image needles in web pages
@@ 3.1 Diverse Long-Context Applications for LCVLMs
To build a long context, we insert the gold passage(s) (the passage with the answer) among a large set of distracting passages retrieved from Wikipedia. For ViQuAE, we use gold passages from KILT [72] since it is constructed upon TriviaQA [73]. For InfoSeek, we choose the lead section of the named entity’s Wikipedia page as the gold reference and remove all examples for which the answer cannot be found in the lead section. Then, we split Wikipedia pages into 100-word passages and incrementally add retrieved passages that do not contain the answer or the named entity as distractors until we reach the given input length $L$. For retrieval, we use the named entity instead of the image itself as the query, because text-based retrieval achieves higher recall and provides harder distractors. In ViQuAE, each example contains a single gold passage, and we insert it at six evenly distributed positions, while in Infoseek, the lead sections often contain hundreds of words, resulting in multiple passages. We randomly shuffle them into three permutations. We use the substring exact match (SubEM) as the metric, following previous work [74]. See more details in Section A.1.
@@
In VH, we obtain the needle images and the target objects from the original dataset. Then, we accompany these needles with multiple negative distractor images until the image haystack reaches a given input length $L$. We report accuracy as the metric following the original work [16]. In MM-NIAH, the haystacks are composed of web documents [75] with interleaved text and images. We include all three tasks of retrieval, counting, and reasoning in our benchmark. Similar to VRAG, we split text in each web document into 100-word passages and add passages and images so that the context achieves the input length. We use SubEM and accuracy to evaluate retrieval and reasoning tasks following the original paper [18]. In the counting task, we use the accuracy of summed needle counts for better robustness. Refer to Section A.2 for more details.
@@
One difference from existing work on many-shot ICL with LCVLMs [79] is that we map the original natural language labels (e.g., food names) to class IDs (e.g., 0, 1, …), requiring models to learn new tasks rather than relying on pre-training knowledge. Section A.3 covers more details.
Summarization (Summ) evaluates an LCVLM’s ability to generate concise outputs from long multimodal documents while preserving all salient information. We choose GovReport [15] (government report summarization) and Multi-LexSum [28] (multi-document legal summarization), as their PDF-formatted documents are long and easily accessible. Our evaluation provides models with PDF-formatted documents rather than OCR-extracted text used in previous works [80, 81]. We truncate the document from the end based on the input length $L$. Following previous work [20], we use LLM-based evaluation for both datasets instead of the commonly used ROUGE-L, as it better reflects human judgment. More details, such as the LLM-based metric, are provided in Section A.4.
Long-document VQA (DocVQA) assesses the model’s aptitude for answering questions that require reasoning over information dispersed across multiple images and text segments within an extended document. We include commonly adopted datasets for evaluating long-document VQA: SlideVQA [82], MMLongBench-Doc [5], and LongDocURL [17]. For documents longer than input length $L$, we truncate the documents evenly from both sides while keeping the answer pages. For shorter documents, we alternately pad the left and right sides with randomly sampled negative documents up to length $L$. However, the padding documents may occasionally contain information related to the question and potentially change the answer. To ensure the validity of questions, we preface each question with the prompt “Based on the Document <Original Doc ID>, answer the following question.” We follow the metrics used in LongDocURL but remove questions with long answers, thereby avoiding LLM-based answer extraction. We list specific details in Section A.5.
@@ 3.3 Standardized Input Length
The input length $L$ is an important factor to consider when we evaluate models’ long-context ability, as longer inputs can provide more information but also challenge models to filter out distracting information. As aforementioned in Section 3.1, we can control the input length $L$ for each dataset either by adjusting the number of passages, images, or exemplars, or by truncating the PDF-formatted documents. This allows us to present each example in our benchmark under multiple standardized input lengths and better understand how performance changes as the context length increases. Specifically, our benchmark provides five input lengths $L$: 8K, 16K, 32K, 64K, and 128K tokens, using binary prefixes $K=2^{10}$, and the input length can be easily extended beyond 128K if needed.
Figure 1: Performance on MMLongBench. We report results for selected frontier models, and the full results of all models are provided in Figure 21. Note that Claude-3.7-Sonnet supports at most 100 images, and we mark the results as N/A for cases with more images (More in Section D.4)
@@ 4 Evaluation and Analysis
With broad task coverage, unified token counting, and standardized input length, we are now able to thoroughly examine LCVLMs’ long-context ability across multiple dimensions. In total, we evaluate 46 LCVLMs on MMLongBench. To the best of our knowledge, our evaluation provides the most thorough and controlled comparison of the vision-language long-context ability on broad real-world applications. These models include closed-source models GPT-4o [10], Claude-3.7 [34], and Gemini 2 and 2.5 [85, 43], as well as open-source model families, such as Qwen2.5-VL [29], InternVL3 [30], and Gemma3 [3]. We also consider position extrapolation methods, such as YaRN [37] and V2PE [52] (See Section D.5). The full list of evaluated models is provided in Table 10. Following existing works [20], we use greedy decoding for all models for consistency and randomly sample 100 examples from each dataset. More details on the experimental setup are in Appendix C.
@@ 4.1 Evaluation on MMLongBench across Tasks and Context Lengths
We present the performance of selected frontier LCVLMs in Figure 1, and the full results of all 46 models are reported in Figure 21. We analyze model performance from multiple perspectives and summarize our main findings as follows:
@@
Reasoning can improve multimodal long-context ability. We include Gemini-2.0-Flash-T in our evaluation, which is the thinking variant of Gemini-2.0-Flash. From the results, we observe that the reasoning ability can consistently improve the Gemini-2.0-Flash on all tasks. While the changes for VRAG, Recall, and ICL are modest, summarization and DocVQA exhibit marked improvements of 25.3% and 10.1%, respectively. Then, Gemini-2.5 models exhibit even stronger performance, which are natively designed as thinking models.
Different models exhibit different strengths. Generally, we find that model performance varies considerably across different tasks. For instance, Qwen2.5-VL-32B outperforms InternVL3-38B on VRAG, but underperforms on NIAH. Similarly, Ovis2-34B excels at summarization but struggles on DocVQA. These findings further support the necessity of a comprehensive benchmark covering diverse downstream tasks. In Sections D.5 and D.6, we also provide additional analysis about the performance of position extrapolations and the lost-in-the-middle phenomenon.
Figure 2: Distribution of long-document VQA (DocVQA) with respect to performance on MM-NIAH variants. We find that the models are concentrated in the coral-shaded areas.
@@ 4.2 Can Needle-in-a-Haystack Tasks Reflect LCVLM’s Overall Long-Context Ability?
Text-image interleaved NIAH tasks are challenging. In Figure 3, we find that even state-of-the-art models like GPT-4o and Gemini-2.5 struggle to surpass 80% accuracy on VH-Multi when the context length is just 8K tokens (approximately 22 images). Most models are just slightly better than a random guess (50%). This demonstrates that locating objects in a large set of images is still hugely challenging for current LCVLMs. See more discussion in Section D.1. Then, we plot the performance of different models on the retrieval, counting, and reasoning tasks of MM-NIAH against their performance on long-document VQA in Figure 2. We find that most models achieve low performance on the counting and reasoning tasks, with scores below 30 and 40, respectively. The difficulty of the tasks and low performance result in poor separability between models. While the retrieval is an easier task, it still does not align well with DocVQA tasks. In short, we find that both VH and MM-NIAH present significant challenges to current LCVLMs, thus showing limited differentiation between models and weak alignment with other tasks.
@@
NIAH tasks fail to reflect overall long-context abilities. As shown in Figure 4, none of the NIAH tasks exhibit strong correlation with the broader set of long-context tasks. This suggests that performance on NIAH tasks may not be a reliable indicator of general long-context capabilities. In particular, Visual Haystack (VH) tasks show especially low correlations due to their high difficulty, as discussed above, which results in limited ability to distinguish between models. In MM-NIAH, counting and reasoning tasks show weak correlations with several downstream tasks, with coefficients below 0.8. The retrieval task also shows weak alignment with ICL performance. Interestingly, simpler tasks — like retrieval with a single needle in unrelated essays — tend to correlate better with diverse task categories, which is consistent with our findings in Figure 2. We further examine the differences between text-based and image-based needles in Section D.2.
Figure 5: Spearman’s $\rho$ between all categories with $L=$128K. For each category, the Avg excludes the correlation with itself.
@@ 4.3 Weak Correlation Across Categories Calls for Diverse Evaluation
We perform a cross-category correlation analysis of model performance. We find that different categories do not consistently show strong correlation ($<0.85$) with each other, as shown in Figure 5. Specifically, VRAG and NIAH closely correlate because retrieval is the central capability of both tasks. A further investigation shows that VRAG achieves its highest correlation (of 0.93) with the retrieval task (MM-NIAH-Ret) in Figure 13, reinforcing the shared emphasis on retrieval. Meanwhile, summarization and long-document VQA show a high correlation of 0.88, likely due to their shared input format — image-based PDF documents. This suggests that image types affect category correlations. In contrast, ICL tasks show relatively weak correlations with other categories. The ICL tasks evaluate models’ ability to induce new classification rules from numerous exemplars, a skill orthogonal to recalling facts in long contexts. This further demonstrates that model developers should consider various long-context skills to draw a more holistic picture of LCVLMs. See Section D.3 for detailed dataset-level correlations and additional category-wise insights.
Long-document VQA is a reliable proxy for long-context capabilities. As shown in Figure 5, long-document VQA achieves the highest average correlation with other categories, indicating that it is more aligned with the broader range of long-context tasks. For example, questions from LongDocURL [17] cover not only simple retrieval but also complex understanding and reasoning. Meanwhile, long-document VQA exhibits the smallest standard deviation, showing that it is also stable and balanced. Taken together, these findings suggest long-document VQA is a more representative and reliable proxy than the commonly adopted NIAH for reflecting overall system performance, allowing model developers to iterate more rapidly without the overhead of full-scale evaluation.

### v3

@@ 1 Introduction
To support such capabilities, researchers have proposed various techniques to extend the context windows of LVLMs, as seen in models such as LongVILA [9] and GPT-4o [10]. However, the development of effective evaluation benchmarks is lagging behind. It remains unclear how well current LCVLMs perform in long-context settings, what types of tasks they struggle with, and how robust they are to input length variation. Here, we take a closer look and find that existing benchmarks suffer from the following shortcomings and provide a summary of key feature comparisons in Table˜1:
@@
Table 1: Comparison of benchmarks for LCVLMs: MM-NIAH [18], Visual Haystack [16], MMNeedle [11], MMLongBench-Doc (MMLB-Doc) [5], M-Longdoc [21], LongDocURL [17], and our MMLongBench. “Summ” and “DocVQA” refer to summarization and long-document visual question answering (VQA). “Mixed” indicates both natural and synthetic images. “LL” denotes the number of input tokens, and “$L$ Control” means counting both text and image tokens together.
@@ 2 Related Work
In contrast, the evaluation of LVLMs’ long-context capability remains limited. Existing benchmarks only involve either NIAH [18, 11, 16, 65] or long-document VQA [5, 17], lacking comprehensive coverage across diverse vision-language applications. As a result, frontier LCVLMs [10, 33] only report long-context performance on other modalities, such as long video [9, 66, 67] or long audio [33, 10], and neglect the prevalent use cases of long-context vision-language inputs. While recent MileBench [68] claims to be a comprehensive long-context benchmark with various text-image tasks, our closer inspection reveals that it actually contains a lot of short-context tasks, and the average length is only about 9K tokens. Datasets like DocVQA [69], WebQA [70], and OCRVQA [71] contain only one image per sample and minimal context, making MileBench unqualified as a true long-context benchmark. In this work, beyond existing long video and audio benchmarks, we introduce the first comprehensive benchmark that evaluates a wide range of vision-language downstream applications.
@@ 3 Our Benchmark — MMLongBench
In our work, we seek to address the limitations of current benchmarks by meeting the following criteria: i) broad coverage of both diverse vision-language downstream tasks and different image types, ii) a unified token-count method across different modalities and datasets, and iii) multiple standardized context lengths for each example, ranging from 8K to 128K tokens. In this section, we describe the task categories and datasets included in MMLongBench, highlighting how they address the limitations of existing evaluation benchmarks. An overview of MMLongBench is provided in Table˜2, and several concrete examples are given in Appendix˜E.
@@
$M$-NIAH-Ret SubEM/Acc Mixed 1,200 Retrieve text/image needles in web pages
$M$-NIAH-Count Acc Mixed 1,178 Count text/image needles in web pages
$M$-NIAH-Reason SubEM/Acc Mixed 1,158 Reason about text/image needles in web pages
@@ 3.1 Diverse Long-Context Applications for LCVLMs
To build a long context, we insert the gold passage(s) (the passage with the answer) among a large set of distracting passages retrieved from Wikipedia. For ViQuAE, we use gold passages from KILT [72] since it is constructed upon TriviaQA [73]. For InfoSeek, we choose the lead section of the named entity’s Wikipedia page as the gold reference and remove all examples for which the answer cannot be found in the lead section. Then, we split Wikipedia pages into 100-word passages and incrementally add retrieved passages that do not contain the answer or the named entity as distractors until we reach the given input length $L$. For retrieval, we use the named entity instead of the image itself as the query, because text-based retrieval achieves higher recall and provides harder distractors. In ViQuAE, each example contains a single gold passage, and we insert it at six evenly distributed positions, while in Infoseek, the lead sections often contain hundreds of words, resulting in multiple passages. We randomly shuffle them into three permutations. We use the substring exact match (SubEM) as the metric, following previous work [74]. See more details in Section˜A.1.
@@
In VH, we obtain the needle images and the target objects from the original dataset. Then, we accompany these needles with multiple negative distractor images until the image haystack reaches a given input length $L$. We report accuracy as the metric following the original work [16]. In MM-NIAH, the haystacks are composed of web documents [75] with interleaved text and images. We include all three tasks of retrieval, counting, and reasoning in our benchmark. Similar to VRAG, we split text in each web document into 100-word passages and add passages and images so that the context achieves the input length. We use SubEM and accuracy to evaluate retrieval and reasoning tasks following the original paper [18]. In the counting task, we use the accuracy of summed needle counts for better robustness. Refer to Section˜A.2 for more details.
@@
One difference from existing work on many-shot ICL with LCVLMs [79] is that we map the original natural language labels (e.g., food names) to class IDs (e.g., 0, 1, …), requiring models to learn new tasks rather than relying on pre-training knowledge. Section˜A.3 covers more details.
Summarization (Summ) evaluates an LCVLM’s ability to generate concise outputs from long multimodal documents while preserving all salient information. We choose GovReport [15] (government report summarization) and Multi-LexSum [28] (multi-document legal summarization), as their PDF-formatted documents are long and easily accessible. Our evaluation provides models with PDF-formatted documents rather than OCR-extracted text used in previous works [80, 81]. We truncate the document from the end based on the input length $L$. Following previous work [20], we use LLM-based evaluation for both datasets instead of the commonly used ROUGE-L, as it better reflects human judgment. More details, such as the LLM-based metric, are provided in Section˜A.4.
Long-document VQA (DocVQA) assesses the model’s aptitude for answering questions that require reasoning over information dispersed across multiple images and text segments within an extended document. We include commonly adopted datasets for evaluating long-document VQA: SlideVQA [82], MMLongBench-Doc [5], and LongDocURL [17]. For documents longer than input length $L$, we truncate the documents evenly from both sides while keeping the answer pages. For shorter documents, we alternately pad the left and right sides with randomly sampled negative documents up to length $L$. However, the padding documents may occasionally contain information related to the question and potentially change the answer. To ensure the validity of questions, we preface each question with the prompt “Based on the Document <Original Doc ID>, answer the following question.” We follow the metrics used in LongDocURL but remove questions with long answers, thereby avoiding LLM-based answer extraction. We list specific details in Section˜A.5.
@@ 3.3 Standardized Input Length
The input length $L$ is an important factor to consider when we evaluate models’ long-context ability, as longer inputs can provide more information but also challenge models to filter out distracting information. As aforementioned in Section˜3.1, we can control the input length $L$ for each dataset either by adjusting the number of passages, images, or exemplars, or by truncating the PDF-formatted documents. This allows us to present each example in our benchmark under multiple standardized input lengths and better understand how performance changes as the context length increases. Specifically, our benchmark provides five input lengths $L$: 8K, 16K, 32K, 64K, and 128K tokens, using binary prefixes $K=2^{10}$, and the input length can be easily extended beyond 128K if needed.
Figure 1: Performance on MMLongBench. We report results for selected frontier models, and the full results of all models are provided in Figure˜23. Note that Claude-3.7-Sonnet supports at most 100 images, and we mark the results as N/A for cases with more images (More in Section˜D.4)
@@ 4 Evaluation and Analysis
With broad task coverage, unified token counting, and standardized input length, we are now able to thoroughly examine LCVLMs’ long-context ability across multiple dimensions. In total, we evaluate 46 LCVLMs on MMLongBench. To the best of our knowledge, our evaluation provides the most thorough and controlled comparison of the vision-language long-context ability on broad real-world applications. These models include closed-source models GPT-4o [10], Claude-3.7 [34], and Gemini 2 and 2.5 [85, 43], as well as open-source model families, such as Qwen2.5-VL [29], InternVL3 [30], and Gemma3 [3]. We also consider position extrapolation methods, such as YaRN [37] and V2PE [52] (See Section˜D.6). The full list of evaluated models is provided in Table˜10. Following existing works [20], we use greedy decoding for all models for consistency and randomly sample 100 examples from each dataset. More details on the experimental setup are in Appendix˜C.
@@ 4.1 Evaluation on MMLongBench across Tasks and Context Lengths
We present the performance of selected frontier LCVLMs in Figure˜1, and the full results of all 46 models are reported in Figure˜23. We analyze model performance from multiple perspectives and summarize our main findings as follows:
@@
Reasoning can improve multimodal long-context ability. We include Gemini-2.0-Flash-T in our evaluation, which is the thinking variant of Gemini-2.0-Flash. From the results, we observe that the reasoning ability can consistently improve the Gemini-2.0-Flash on all tasks. While the changes for VRAG, Recall, and ICL are modest, summarization and DocVQA exhibit marked improvements of 25.3% and 10.1%, respectively. Then, Gemini-2.5 models exhibit even stronger performance, which are natively designed as thinking models. See more results with newly added models in Section˜D.5.
Different models exhibit different strengths. Generally, we find that model performance varies considerably across different tasks. For instance, Qwen2.5-VL-32B outperforms InternVL3-38B on VRAG, but underperforms on NIAH. Similarly, Ovis2-34B excels at summarization but struggles on DocVQA. These findings further support the necessity of a comprehensive benchmark covering diverse downstream tasks. In Sections˜D.6 and D.7, we also provide additional analysis about the performance of position extrapolations and the lost-in-the-middle phenomenon.
Figure 2: Distribution of long-document VQA (DocVQA) with respect to performance on $M$-NIAH variants. We find that the models are concentrated in the coral-shaded areas.
@@ 4.2 Can Needle-in-a-Haystack Tasks Reflect LCVLM’s Overall Long-Context Ability?
Text-image interleaved NIAH tasks are challenging. In Figure˜3, we find that even state-of-the-art models like GPT-4o and Gemini-2.5 struggle to surpass 80% accuracy on VH-Multi when the context length is just 8K tokens (approximately 22 images). Most models are just slightly better than a random guess (50%). This demonstrates that locating objects in a large set of images is still hugely challenging for current LCVLMs. See more discussion in Section˜D.1. Then, we plot the performance of different models on the retrieval, counting, and reasoning tasks of MM-NIAH against their performance on long-document VQA in Figure˜2. We find that most models achieve low performance on the counting and reasoning tasks, with scores below 30 and 40, respectively. The difficulty of the tasks and low performance result in poor separability between models. While the retrieval is an easier task, it still does not align well with DocVQA tasks. In short, we find that both VH and MM-NIAH present significant challenges to current LCVLMs, thus showing limited differentiation between models and weak alignment with other tasks.
@@
NIAH tasks fail to reflect overall long-context abilities. As shown in Figure˜4, none of the NIAH tasks exhibit strong correlation with the broader set of long-context tasks. This suggests that performance on NIAH tasks may not be a reliable indicator of general long-context capabilities. In particular, Visual Haystack (VH) tasks show especially low correlations due to their high difficulty, as discussed above, which results in limited ability to distinguish between models. In MM-NIAH, counting and reasoning tasks show weak correlations with several downstream tasks, with coefficients below 0.8. The retrieval task also shows weak alignment with ICL performance. Interestingly, simpler tasks — like retrieval with a single needle in unrelated essays — tend to correlate better with diverse task categories, which is consistent with our findings in Figure˜2. We further examine the differences between text-based and image-based needles in Section˜D.2.
Figure 5: Spearman’s $\rho$ between all categories with L=L=128K. For each category, the Avg excludes the correlation with itself.
@@ 4.3 Weak Correlation Across Categories Calls for Diverse Evaluation
We perform a cross-category correlation analysis of model performance. We find that different categories do not consistently show strong correlation (<0.85<0.85) with each other, as shown in Figure˜5. Specifically, VRAG and NIAH closely correlate because retrieval is the central capability of both tasks. A further investigation shows that VRAG achieves its highest correlation (of 0.93) with the retrieval task (MM-NIAH-Ret) in Figure˜15, reinforcing the shared emphasis on retrieval. Meanwhile, summarization and long-document VQA show a high correlation of 0.88, likely due to their shared input format — image-based PDF documents. This suggests that image types affect category correlations. In contrast, ICL tasks show relatively weak correlations with other categories. The ICL tasks evaluate models’ ability to induce new classification rules from numerous exemplars, a skill orthogonal to recalling facts in long contexts. This further demonstrates that model developers should consider various long-context skills to draw a more holistic picture of LCVLMs. See Section˜D.3 for detailed dataset-level correlations and additional category-wise insights.
Long-document VQA is a reliable proxy for long-context capabilities. As shown in Figure˜5, long-document VQA achieves the highest average correlation with other categories, indicating that it is more aligned with the broader range of long-context tasks. For example, questions from LongDocURL [17] cover not only simple retrieval but also complex understanding and reasoning. Meanwhile, long-document VQA exhibits the smallest standard deviation, showing that it is also stable and balanced. Taken together, these findings suggest long-document VQA is a more representative and reliable proxy than the commonly adopted NIAH for reflecting overall system performance, allowing model developers to iterate more rapidly without the overhead of full-scale evaluation.